OpenStudy (anonymous):

A pharmacist wants to mix an ointment that is 6% zinc oxide with an ointment that is 12% zinc oxide to make 30 grams of an ointment that is 10% zinc oxide. How many grams of each ointment should the pharmacist mix together? Can anyone breakdown this problem for me, so that I can learn how to do these types of problems. Thanks

OpenStudy (anonymous):

Mixture problems are hard. Take it a step at a time. We know that the end product is .10 x (30) The first ingredient is .06 time some amount x, the second ingredient is .12 x (30 - x) if we put it all together \[.06x + .12(30 - x) = .10(30)\] \[.06x + 3.6 - .12x = 3.0\] \[-.06x + 3.6 = 3.0\] \[-.06x = -0.6\] \[x = 10\] so we need 10 grams of the 6% and 20 grams of the 12%

OpenStudy (anonymous):

Thank you so much for helping me with my problem.